### v1

# Application of AI to formal methods — an analysis of current trends

Abstract
With artificial intelligence (AI) being well established within the daily lives of research communities, we turn our gaze toward an application area that appears intuitively unsuited for probabilistic decision-making: the area of formal methods (FM). FM aim to provide sound and understandable reasoning about problems in computer science, which seemingly collides with the black-box nature that inhibits many AI approaches. However, many researchers have crossed this gap and applied AI techniques to enhance FM approaches. As this dichotomy of FM and AI sparked our interest, we conducted a systematic mapping study to map the current landscape of research publications. In this study, we investigate the previous five years of applied AI to FM (2019–2023), as these correspond to periods of high activity. This investigation results in 189 entries, which we explore in more detail to find current trends, highlight research gaps, and give suggestions for future research.
keywords:
formal methods , artificial intelligence , machine learning , systematic mapping study
journal: Information and Software Technology
[…]
[1]organization=Johannes Kepler University Linz, organization=Institute of Software System Engineering, city=Linz, postcode=4040, country=Austria
[…]
[2]organization=Heinrich Heine University Düsseldorf, organization=Faculty of Mathematics and Natural Sciences, organization=Institute of Computer Science, city=Düsseldorf, postcode=40225, country=Germany
{highlights}
Systematic mapping of AI application within the formal methods domain.
Analyzed tally of 189 primary studies from the last five years (2019–2023).
Data set of all 457 found publications, including the mapped tally of 189 processed studies, made publicly available at https://github.com/hhu-stups/ai4fm-studies.
Formulation of research questions to guide future research.
1 Introduction
Artificial intelligence (AI), and especially the subdomain of machine learning (ML), is becoming increasingly relevant for all industries and increasingly penetrates research communities, as several studies show [1, 2, 3, 4]. Consequently, AI applications also gained popularity in software development  [5, 6].
In this study, we ask whether this trend is also observable within the formal methods communities. Formal methods (FM) correspond to a mathematically rigorous approach to software and systems development in which the concern is to provide a formal representation of a hardware, software, or systems engineering problem [7]. This representation then lends itself to mathematical assessments such as correctness, soundness, or well-definedness. Furthermore, these assessments can be easily reproduced and followed by the practitioner. The needed guarantees and the rigorousness of these methods seem to stand in tension with AI applications, which often show nondeterministic behavior and mostly correspond to black-box techniques. However, much like in software development, AI-driven tools might still supplement the formal development process without impeding the correctness of the results. And from personal experience, we noticed an increase in publications applying AI techniques to FM in recent years, suggesting that the FM field is following the overall trend as well. To substantiate this and to obtain an overview of which FM subdomains already utilize AI, we performed this systematic mapping study (SMS) [8]. The primary goal is to observe the development of AI applications quantitatively to the field of FM.
In this work, we report on our journey and the results of our SMS, accessing the quantitative nature of the topic of AI application in the FM domain. This represents a typical goal for this kind of study, as pointed out by Kitchenham et al. [9]. Mainly, we are interested in which FM domains are targeted by AI improvement already, which AI or ML applications have been investigated in the literature so far, where potential research gaps exist, and whether there appears to be consensus within the FM domains of which AI techniques seem most beneficial. A challenge to this assessment is the wide variety of topics on both sides. AI and FM are huge topics, and their potential overlap may be considerable. This study aims to be the first assessment of the general research landscape, i.e., we assume a mainly quantitative lens on the topic. Consequently, we highlight the areas which received more attention in the literature as well as those that received less attention. We refrain, however, from doing a qualitative assessment (e.g., a systematic literature review) our goal is to gain an initial overview of the field’s landscape.
While our search process resulted in a data set of 457 highly relevant publications covering more than 50 years of research applying AI to FM, we investigated only articles from the last five years in more detail. This period is chosen as it corresponds to a peak in the number of publications as indicated by previously mentioned studies and also allows us to focus on the most recent developments. Consequently, we end up with a tally of 191 primary studies, two of which are pure data set presentations. We analyzed the remaining 189 studies to determine which AI techniques were found to be applicable in which FM areas. Further, we derive research suggestions from our insights by pointing out research gaps that might further mature the field. Both the data set of the last five years as well as the complete data set of 457 publications are made publicly available to interested researchers at https://github.com/hhu-stups/ai4fm-studies.
The rest of this mapping study is structured as follows: Section 2 introduces the relevant subdomains of FM and AI. It will also provide an overview of our understanding of AI. We present our leading research questions in Section 3 and detail our search and selection process in Section 4. Section 5 presents the results of the search while Section 6 engages in a deeper discussion about these results. In Section 7, we discuss potential threats to the validity of our conducted systematic mapping study. Section 8 presents related studies we found during our research and compares them against our efforts. Finally, we conclude in Section 9.
2 Background
2.1 Systematic mapping studies
Systematic mapping studies, as proposed by Kitchenham et al. [10], aim to assess the quantitative nature of a research field. Here, the goal is to give an overview of the available research. SMS complement systematic literature reviews by comparing research contributions and their qualitative impact.
SMS are typically performed using a multi-step approach. First, the field of interest is selected. Then, research questions are formulated, and inclusion criteria (IC) and exclusion criteria (EC) are defined. IC and EC will serve as filters to help select relevant results.
After that, the search query is crafted. The goal is to apply the search query to the scientific meta-search engines to retrieve a good corpus of contributions that answer the research questions. With the retrieved corpus, the IC and EC are applied. If indicated, snowballing is done, usually until a stable closure of the corpus is reached. With a corpus constructed, the aim is to answer the research questions.
2.2 Formal methods
Formal methods (FM) describe a mathematically rigorous approach to design software as well as hardware systems, and are concerned with analysis, validation, and verification at any part of the respective system’s life cycle [11, 12]. Employing FM allows to formulate precise statements of desired functionality or requirements in form of a formal model (or formal specification) while not constraining the possible implementation thereof. The formal model is represented in a mathematically approachable form which lends itself to reasoning and hence allows rigorous analysis of critical properties such as correctness, safety, soundness, or well-definedness.
In the following, we overview FM creation and reasoning methods most relevant to this works.
Model checking.
Model checking [13, 14] explores a program’s or systems’ state space. The state space sets all possible value constellations the program can achieve. In its most basic form, explicit state model checking, model checking aims to explore all reachable states and check if they are faulty, i.e., if they violate any specified properties. If this is the case, a counter-example is found. Otherwise, the system works correctly, as no faulty states are reachable.
There are many flavors besides explicit state model checking. Noteworthy for this study are linear temporal logic (LTL) model checking, symbolic model checking, and statistical model checking. LTL model checking encodes temporal properties of interest into a formula that is checked against execution traces of the state space. Symbolic model checking abstracts the state space and makes a symbolic evaluation of a complex expression, resulting in the ability to handle even infinite state spaces for the price of precision. Statistical model checking investigates state spaces created by assigning probabilities on transitions between states and provide probabilistic guarantees.
Theorem proving.
A core application of theorem proving (TP) is to evaluate statements about a formal model’s internal consistency and behavioral integrity. When talking about theorem proving, we usually assume some formal representation of the problem for which proofs are formulated and discharged. Discharging proofs can either be done by a fully automatic theorem prover (ATP) [15] that finds a solution on its own, or an interactive theorem prover (ITP) [16] that requires human input for individual proof steps and transformations. ITPs provide a user interface to keep track of progress, sub-goals, and available properties.
SAT/SMT solving.
The Boolean satisfiability problem (SAT) [17] refers to finding the right model for a boolean formula to satisfy the formula. Historically, SAT was the first problem found to be NP-complete and has since drawn a broad research interest.
Satisfiability modulo theories (SMT) [18] is a more generalized form of the SAT which enriches the problem setting with various theories such as arithmetics, data structures, or set theory.
In the context of FM, SAT and SMT solving find applications in theorem proving [19], bounded model checking [20], equivalence checking [21] or test generation [22].
Synthesis.
Under the term of synthesis, we group all attempts automatically or semi-automatically create (parts of) formal models during the formal development process, such as the generation of models from natural language specification or vice versa. We further group the idea of automated model repair under this aspect, where for a model with some violation, a fix is synthesized that rectifies the violation.
Other categories.
In the scope of this mapping study, we may encounter FM approaches and techniques, which do not belong to any previous group. We will group these as other FM techniques. For instance, this includes meta-approaches such as selecting the right verification algorithm, general program analysis, or termination analysis.
2.3 Artificial intelligence
Artificial Intelligence (AI) is a catch-all phrase for various techniques and approaches that aim to imitate intelligent decision-making. Simmons and Chappell [23] define the term as “[…] behavior of a machine which, if a human behaves in the same way, is considered intelligent”. They further emphasize the problem with the term itself, pointing out the difficulty of precisely defining what it means for a human to be intelligent. Nevertheless, AI is regularly used in daily life and the research community.
An important AI paradigm is machine learning (ML). The goal of ML algorithms is to observe and extract patterns within the data which can then be applied to the automation of a given task  [24]. Instead of manually defining a set of rules for the pattern recognition, machine learning algorithms uncover such rules in an automated fashion, a process referred to as learning. ML algorithms experience the given data in some manner and improve their exhibited performance to solve the task at hand gradually, the more experience they gather.
Below, we list and briefly explain AI and ML algorithms which are most relevant to this mapping study. For the reader familiar with AI, the selection seems mostly ML rather than AI. Indeed, we aimed to find applications of AI to FM, and the search query we will discuss later was prepared accordingly. However, in order to bring a result up front, most contributions use some form of ML.
Neural networks and deep learning.
Neural networks (NNs) [25] build the foundation of what is known today as the area of deep learning [26]. A deep neural network (DNN) consists of a layer of inputs, a layer of outputs, and one or more hidden layers. Inputs are forwarded layer-wise and combined first by a weighted sum and a follow-up non-linear transformation. The clue is the training of a DNN, which starts the network with randomly initialized weights for the summations above but adjusts them gradually in the learning process to converge to the desired function.
A considerable benefit of neural networks lies in their variety. Different approaches and architectures exist for deep learning in specialized environments with differently shaped data. The most important for this work are convolutional neural networks (CNNs) [27] for images, recurrent neural networks (RNNs) [28] for sequential data, the transformer architecture [29] for language processing, and graph neural networks (GNNs) [30]. Today, deep learning is considered one of the most popular topics in ML [31].
Reinforcement learning.
Reinforcement learning (RL) [32] is foremost a machine learning paradigm. Given an environment, a set of actions that change the environment, and a reward function that quantifies a given environmental state, the RL agent’s goal is to learn a policy over their available actions which maximizes the cumulative reward over time. Initially, the agent does not know the environment or the task it is supposed to solve but can only learn from feedback through the received rewards [33].
Natural language processing.
Natural language processing (NLP) [34] describes the research area of, as the name suggests, processing and evaluating texts of human language. This includes information extraction, language translation tasks, text classification, semantic analysis, and natural language generation and dialogue systems [34, 35].
One of the most recent developments is large language models (LLMs) [36] which gained significant popularity and even worldwide attention with the release of ChatGPT [37].
Genetic/evolutionary algorithms.
Genetic and evolutionary algorithms (EAs) [38] describe a family of optimization algorithms that are rooted in the phenomenon of evolution. An EA starts with a randomly assembled population of candidates for a given optimization problem and a fitness function. The fitness function measures how well a candidate solves the problem. Parent candidates are chosen via a random selection process that is nonetheless influenced by each individual’s fitness. These are adapted using mutation or recombination to produce a new child population. Repeating this process converges the individuals toward suitable solutions for the problem at hand.
Statistical approaches.
In contrast to the abovementioned techniques, we consider techniques from more classical, statistical ML. These differ from the above methods as they do not depend on deep learning and are typically faster to compute in comparison. This does, however, not imply that they are not competitive. The considered statistical approaches include support vector machines (SVM) [39], logistic and linear regression (LR) [40], k nearest neighbors (KNN) [41], support vector machines (SVM) [42], decision trees (DT) [43] and random forests (RF) [44], or Bayesian inference approaches [45].
2.4 Contribution classification
We classified the contributions that we selected via the PICO [10] methodology. Thus, we facilitate a clear differentiation between them and allow for a more refined analysis.
The different types of contributions are distinguished as follows:
Tool.
Contributions that provide practical means to solve an FM problem, for instance, in the form of a binary or source code release.
Tool enhancement.
Contributions that aim to enhance or complement an existing tool.
Approach.
Contributions that propose an overall style or idea to overcome a problem. An approach is a generalized concept that describes how to react or overcome a problem in depth. The approach remains more theoretical and does not involve tested or empirically proven steps.
Methodology.
Contributions that realize and test an approach.
Framework.
Contributions that propose a conceptual structure intended to support or guide the construction or expansion of a solution for an actual problem.
Case study.
Contributions applying an approach, a methodology, or a framework to a real-world problem.
Benchmark.
Contributions that test an existing tool or methodology.
Idea.
Contributions that very briefly overview how a problem may be solved. Compared to an approach, it does not lay out larger theories but serves as a quick pitch.
Training data.
An openly accessible data set that claims to be reusable outside of replication of the contribution it was originally used in.
We further divide the publication venues into the following categories:
Workshop and short papers.
We grouped contributions in this category if they are (a) rather short (4 pages or less) or (b) if the venue they appeared in has workshop character, i.e., is classified as a workshop, calls itself a workshop, or focuses on the in-person presentation of results but requires submission of an extended abstract.
Conference proceedings.
Here, we group all the contributions that are published in the proceedings of a conference. From a full conference paper, we expect a more in-depth explanation of the result than compared to a workshop or short paper, ultimately resulting in an overall more sizable contribution.
Journal articles.
Here, we group all contributions published as part of a journal volume. As journal articles tend to consist of more pages and go through a more thorough reviewing process, the results of journal articles are expected to have a broader scientific basis for their claims.
Book chapters.
As the name suggests, these contributions are part of a larger collection of articles bundled in one book. As the production of scientific books takes a significant amount of time, the reader expects a high quality of the provided scientific contribution that is considered, by that time, state of the art.
3 Goal & Research questions
To get an overview of the field, we divide our research questions (RQs) into two main areas. First, RQs 1.1, 1.2 and 1.3 focus on the overall demographics of the research field. With these questions, we want a basic quantitative overview to explore the field’s maturity. Second, RQs 2.1, 2.2, 2.3 and 2.4 focus on the content of the contributions and aim to map out the field, highlighting potential gaps and showcasing quantitative interest in respective sub-topics. Our research questions are defined as follows.
  1. RQ 1:
    Demographics of the research area
    1. 1.1
      What is the research publication timeline, and is there a trend?
      Rational: We are interested to know whether the field experiences a growth or decline in interest, or if no trends are detectable at all.
    2. 1.2
      Which publication venues are most frequent in the field?
      Rational: We want to explore the field’s maturity, whereby books, book chapters, and journal articles indicate more mature results.
    3. 1.3
      What are the provided main contributions by the primary studies?
      Rational: We want to learn where the community focuses its efforts, e.g., more on the theoretical groundwork or more on practical application.
  2. RQ 2:
    Content type of contributions
    1. 2.1
      Which AI techniques and tools were used? Are there any prevalent choices within the community?
      Rational: We want to know if there is an indication that some AI techniques are more relevant (or at least popular) than others.
    2. 2.2
      What are the application areas of AI in FM? Are there any commonly found FM techniques or tools? Rational: We are curious about which FM techniques currently get the most attention or if there are dark spots in the literature.
    3. 2.3
      What is the distribution of AI types in the different FM application areas?
      Rational: Here we want to overlap the results of RQs 2.1 and $2.$
    4. 2.4
      Are the studies’ employed data sets publicly available? Rational: Especially in machine learning, having the source material available and accessible for experiments benefits the scientific community.
4 Search strategy
In the following, we explain our approach for aggregating the contributions for this mapping study which we have visualized in Fig. 1. For this, we follow the suggested multi-step approach by Petersen et al. [8], which consists of an initial search in meta-search engines, followed up by a snowballing process. As it will show, our selection of topics provided a challenging environment which has overlapping search terms with related research areas. To avoid a result set of tens of thousands of potential studies, we tweaked our search strategy to deal with these challenges.
A vital role played the IC & EC, which we applied twice. The first application was made after the initial search to remove unrelated contributions that would introduce an overheat into the snowballing. The second time was again after the snowballing to filter wrongly selected contributions.
Figure 1: Visualization of the search process
4.1 Employed search query
Finding a search query that would return a manageable amount of relevant contributions proved difficult. The main problem was harnessing the bandwidth of terms used in the AI and FM domains. Hereby, the difficulty lies in the often ambiguous use of terms like “formal method”, which can be used in the sense of formal methods as outlined in Section 2.2 or as a term to describe an only somewhat formal approach to a problem. The situation is even more difficult for AI as AI and ML are often used interchangeably. In contrast, essential publications may only use the specific name of an applied technique without naming the fields of AI or ML.
Another issue was the cross-pollution of our results with studies from the related research field of verification for AI and ML systems, which is essentially the other direction in which FM is applied to AI. This field is also a highly relevant topic and shares the majority of potential search key phrases, further inflating our results tally.
As a reaction, we experimented with different approaches and probed the quantity and quality of their respective results to find a suitable search query. Highly abstract queries such as
“formal methods & artificial intelligence”
produce multiple thousands of results which we deemed impossible to process in a reasonable amount of time and which contained many studies outside of our intended scope. From probing some results as well as personal experience, we also knew that many relevant publications tend not to use these high-level terms, but use nomenclature specific to their sub-community instead. For instance, research regarding the formal B method tends to simply use “B method” instead of “formal method” as keyword. On the other hand, we noticed how relevant publications seemed also to prefer to use the concrete names of the applied AI algorithms rather than stating they used AI in a superficial manner. Therefore, we were concerned that a search query that is too abstract would not be able to penetrate the relevant fields well enough.
As a result, we settled on a more complex query containing the precise terminology of specific algorithms. For this, we distinguish between two sets of terms, the AI-terms (Fig. 2(b)) and the FM-terms (Fig. 2(a)). Both correspond to a disjunction of a set of selected terms from the respective field. The terms themselves were selected based on previous experiences in the field following PICO Kitchenham et al. [10].
The final search query (Fig. 2(c)) now makes use of both term aggregates to produce a corpus of primary studies in a strict and targeted manner. That is, we only looked for studies that make use of at least one AI term and at least one FM term in their titles, in their abstracts, and in their keywords. The reasoning behind this is to get a good initial penetration of all key fields while relevant but missed entries are found in the later search stages via snowballing.
FM-terms := “formal method” OR “formal model” OR “specification” OR “rigorous” OR “prove*” OR “solv*” OR “automated reason*” OR “formal synth*” OR “model check*” OR “model repair” OR “premise selec*” OR “SAT” OR “SMT” OR “smtlib”
(a) FM keywords used in the query
AI-terms := “artificial intelligence” OR “AI” OR “machine learn*” OR “supervised learn*” OR “unsupervised learn*” OR “classification” OR “regression” OR “portfolio” OR “deep learn*” OR “neural net*” OR “bayesian net*” OR “reinforcement learn*” OR “reinforcement agent” OR “multi-armed bandit” OR “knn” OR “k nearest neighbours” OR “support vector” OR svm OR “decision tree” OR “random forest” OR “gradient boost*” OR “xgboost” OR “logistic regress*” OR “linear regress*” OR “cluster*” OR “computer vision” OR “nlp” OR “natural language proces*” OR “genetic prog*” OR “genetic algorithm” OR “expert system” OR “inductive logic progr*”
(b) AI keywords used in the query
TITLE(AI-terms) AND TITLE(FM-terms) AND ABSTRACT(AI-terms) AND ABSTRACT(FM-terms) AND KEYWORDS(AI-terms) AND KEYWORDS(FM-terms)
(c) Unified search query
Figure 2: Construction of the used search query. The query consists of a disjunction of search terms for AI and FM, respectively. The unified query (2(c)) enforces that at least one keyword from each subfield is present in the title, the abstract, and the keywords of a respective publication. An asterisk (*) indicates a wildcard character that can match with any sub-word. For instance, “solv*” matches with solver but also with solving.
4.2 Database search and processing
The search was conducted in the last quarter of 2023. We used four meta-search engines [^1] Springer’s engine was not used because by the time of the search it performed extremely poor in dealing with large search queries, large result sets and filtering. as suggested by Petersen et al. [8]: IEEE [^2] https://ieeexplore.ieee.org/Xplore/home.jsp, Scopus  [^3] https://www.scopus.com/search/form.uri?display=basic#basic, ACM [^4] https://dl.acm.org/ and Web of Science (WoS) [^5] https://www.webofknowledge.com/. We used the Guide to Computing Literature for ACM. Furthermore, as the ACM search engine limits the number of wild cards, we split the search terms into subsets, performed the sub-searches, and merged them into the required superset. For WoS, results seemed to differ depending on which institution has access. Therefore, we decided to take the more extensive result set. As far as the search engines allowed, we applied the EC, such as the area of publishing and the publication language. The result size after this step was 1492 entries. This number and all later numbers are without duplicates.
4.3 Inclusion and exclusion criteria
For the 1492 resulting contributions, we made two observations. First, the amount was too large for a snowballing procedure to be feasible. Second, while briefly looking at the corpus, we discovered many contributions that should not have been selected, i.e., contributions that included the term “specification” but meant it in a purely requirement engineering-minded context. Therefore, we decided to apply our IC and EC prematurely to the corpus as far as they applied, i.e., to title and abstract. For this, we read the titles of our studies, and if we could not decide whether the contributions should be included, we also conducted the abstracts, and very rarely, the contributions itself was skimmed.
To achieve this for such a large corpus, the first and second authors passed over all entries individually and decided whether they should be included or excluded. In agreement cases, the respective studies were kept or discarded, respectively. For cases of disagreement, both authors discussed each instance together to find an agreement. The disagreement could not be settled this way in two cases, and the third author was consulted as a tiebreaker. This procedure brought the number of primary studies down to 89 but took a significant amount of time.
The applied IC and EC are detailed in the following.
  1. IC 1:
    The contribution focuses on applying techniques from the domain of AI to the domain of FM.
  2. IC 2:
    The contribution is in English.
  3. IC 3:
    The contribution has at least two pages of content (excluding references).
  4. IC 4:
    If the contribution was submitted in different versions, we took longer. For instance, the journal version was taken for a conference contribution and subsequently published as an extended version. The journal version was taken for a journal-first submission and an invited version.
  5. IC 5:
    The contribution is a data set used for ML.
  1. EC 1:
    The contribution appears to be not reviewed in any form.
  2. EC 2:
    The contribution is not available.
  3. EC 3:
    The contribution not posed in computer science.
  4. EC 4:
    The contribution is unclear (even after reviewing the full text).
  5. EC 5:
    The contribution solely focuses on solving a mathematical optimization problem.
  6. EC 6:
    (During initial skim) The contribution focuses only on common or general constrain-solving techniques (without specializing in SMT/SAT)
  7. EC 7:
    (During initial skim) The contribution focuses on portfolios for SAT solvers without explicitly mentioning any AI/ML technology
  8. EC 8:
    (During initial skim) The contribution is not a primary study, i.e., it is another mapping study, survey, etc.
  9. EC 9:
    (Snowballing) The contribution is concerned with 2Sat, which is solvable in polynomial time.
  10. EC 10:
    (Snowballing) If the contribution proposes a way to synthesize some automaton (e.g., a Markov process), then the produced result must explicitly be used within the context of FM. This means that the synthesized automaton and its analysis with FM’s help are described within the contribution.
For the EC, some additional criteria were added in later steps in response to contributions encountered that were of low value for this study. For instance, EC 6 was added because many contributions focused on constraint solving, thus being closer to mathematics. EC 7 was added in response to a large influx of SAT contributions that mentioned learning but focused on remembering already-seen clauses. EC 8 is a rule that was introduced to not have secondary studies in the result set but to use those studies to ensure that even with a restrictive query, we cover as much of the field as possible. EC 10 was added in the fifth snowballing round, where we got many contributions that synthesized automatons, but no further analysis was mentioned. EC 9 was added as we found 2Sat as a problem not relevant for this study, as solutions are achievable in polynomial time already.
4.4 Snowballing
For the 89 entries, we conducted an extensive snowballing procedure [46] where we conducted both forward and backward snowballing. Five iterations were needed until we eventually reached a closure. This extensive snowballing approach was meant to complement the strict search query and to uncover relevant yet initially missed studies or even subfields of research. Therefore, snowballing served as a means to alleviate these threats to validity.
For the backward snowballing, we consulted the references of a given contribution and selected promising titles or publications that were explicitly highlighted in the respective related work sections. If a publication appeared promising but the contribution was not clear from its title alone, the abstract was consulted as well. For the forward snowballing, we used Google Scholar and applied the same title and abstract procedure. This brought the number of contributions up to 540. After concluding the snowballing, we applied the IC and EC a second time, reducing the amount of contributions to 457.
4.5 Filtering by years
While the 457 results are all highly relevant, the number of contributions was too large to conduct a deeper investigation necessary to answer all the RQs that require a more in-depth consideration of the contributions, i.e., RQs 1.2, 1.3 and 2. Therefore, we restricted the result tally to studies published between 2019 and 2023. While this shifts our focus on the most recent developments in the field, it aligns with the peak in this period we discussed back in Section 1. By only considering studies from 2019 to 2023, we reduced the corpus to 191 primary studies. Of these, 2 studies purely provide data sets for future AI applications in the field of FM without showcasing any form of direct AI application themselves (IC 5). That leaves us with a final tally of 189 primary studies to conduct this mapping study.
5 Results
In the following section, we will discuss the results. For this, we will answer the individual research questions and aim to make a cross-analysis between individual research questions, thus extracting as much information as possible from the existing data set. We refer to Section 6 for a discussion of the observed results.
5.1 RQ 1: Demographics of the research area
5.1.1 RQ 1.1: What is the research publication timeline, and is there a trend?
The whole corpus of 457 primary studies we found was published in the years 1972–2023. Figure 3 displays the histogram of publications per year (see also Appendix A) and highlights that AI applications on FM appear to mirror the trend of general AI application mentioned in Section 1. That is, we can observe a steadily growing interest in the field, especially since around 2005.
Focusing on the last five years only (recall Section 4.5), there still appears to be growing interest overall, while not each year surpasses the respective previous one. For 2019, we have 32, for 2020, we have 48, for 2021, we have 39, for 2022, we have 51, and for 2023, we have 21 contributions. Noteworthy are two observations: First, for the last five years alone, there has been no strict upward trend in the number of publications, We can observe a decline from 2020 to 2021. Second, there were few publications in 2023. We explain this because we searched in the last quarter of 2023 when not all possible studies were published. Given that we do not yet have complete data for 2023, it is impossible to forecast whether the overall trend will continue to grow or if the current peak from 2022 marks a global maximum.
Figure 3: Distribution of contributions over years
5.1.2 RQ 1.2: Which publication venues are most frequent in the field?
Figure 4 (see also Appendix A) displays the publications over the years divided by publication venues. The majority of studies were published as conference papers while journal papers seem to be on the decline again since their peak in 2018.
Within our five-year observation range, 122 contributions were conference papers, followed by 39 journal articles 29 workshop or short papers, and only one (1) book chapter with no full books.
Figure 4: Distribution of publication venues
5.1.3 RQ 1.3: What are the provided main contributions by the primary studies?
Following the definitions of contribution types from Section 2.4, we see the division of the studies into these contribution types in Fig. 5(a) (see also Appendix A). We can see that the majority of contributions were methodologies (118), i.e., practical application of AI to FM, followed by studies with multiple types of contributions (29). Third place was tool contributions (14), followed by improvements for tools (10), followed by frameworks (6), training data (5), case studies (2), benchmarks (2), approaches (2), and finally one (1) idea contribution. Studies with multiple contribution types most commonly provide a methodology (20/29), training data (14/29), or a tool (11/29). See Fig. 5(b) for a complete breakdown.
(a) Singular contribution types
(b) Division of studies’ multiple contributions
Figure 5: Distribution of contribution types of collected primary studies.
5.1.4 Intersecting RQs 1.2 and 1.3
We can gain some additional insights by investigating the individual results in context to each other. Figure 6 (see also Appendix A) shows how the publication venues and the contribution types correlate to each other. We can see that for any contribution type, a heavy focus lies on conference papers. Interestingly, we can see that case studies are only conducted within the scope of journal papers and book chapters. Furthermore, workshops and short papers are primarily used to introduce new methodologies, which seem unexpected, as usually, due to limited space, authors aim to restrict themselves to outlining an approach or an idea.
Figure 6: Contribution type and publication venues. Studies with multiple contribution types are considered individually for each distinct contribution type.
5.2 RQ 2: Content type of contributions
5.2.1 RQ 2.1: Which AI techniques and tools were used?
In Fig. 7(a) (see also Appendix B), we can see that the majority of contributions use NN (70) followed by RL (32). 27 contributions use multiple techniques, while 16 use NLP methods and 14 use EA. A total of 8 studies presented custom algorithms. The rest of the contributions are divided between utilizing decision trees (7), clustering (3), Bayesian inference (3), KNN (2), random forests (2), data mining approaches (2), automaton learning (2), and SVM (1).
Studies that employed multiple algorithms mostly did so in a contrasting manner, i.e., conducted training on numerous different models to see which one performed best for their respective application. A detailed view of the category of multiple algorithms is given in Fig. 7(b) (see also Appendix B). Here, we can see that if multiple techniques are used, they mainly rely on NN (13/31), random forest (12/31), and some variant of tree learners (12/31). Interestingly, SVM (6/31) and KNN (5/31) applications are more predominant in a setting with multiple algorithms compared to being the sole focus of a study. We explain this by KNN or SVM being simple enough algorithms for authors to train them in a small amount of time. Further, they do need less fine-tuning than, say, NNs, lending themselves as baseline approaches for comparisons.
(a) Quantity of the AI types in contributions
(b) Detail view for the category “Mult. algorithms” of Fig. 7(a)
Figure 7: Overview of AI contributions. Underrepresented AI types are aggregated as Other with a number in parenthesis indicating how many sub-techniques were included.
Intersecting RQs 2.1 and 1.3
Figure 8 (see also Appendix B) shows the intersection of RQs 2.1 and 1.3. The wide focus on methodologies was already assessed with Fig. 5(a). In Fig. 8, we can see that the amount of NN contributions remained focused on methodologies (37/189) or as part of a broader application of techniques (13/189) or a tool (8/189).
Figure 8: Distribution of contribution type and AI techniques
Intersecting RQs 2.1 and 1.1
Figure 9 (see also Appendix B) show the intersection of RQs 2.1 and 1.1. Here, we can see that the amount of NN contributions remained steady over the years, similar to the amount of RL contributions. Noteworthy is that there is no notable decline in any technique over the observed period. The alternating nature of the overall amount of contributions was already found back in Section 5.1.
Figure 9: Distribution of year and AI techniques
5.2.2 RQ 2.2: What are the application areas of AI in FM?
Figure 10(a) (see also Appendix B) gives an overview of the FM techniques found in the contributions. TP is leading by a large quantity with 85 entries. SAT is next with 45 entries. Synthesis approaches come in third place with 19 entries, followed by model checking (18) and SMT solving (13). The remaining entries fall into algorithm selection, program analysis, and termination analysis.
In Fig. 10(b) (see also Appendix B), we can see a more detailed investigation of the topic of theorem proving. Here, we can see that premise, axiom, and clause selection received the most attention (27/85), followed by proof search (21/85) and proof synthesis (18/85). After that, a significant gap exists in a row of smaller topics.
In Fig. 10(c), we see the division of SAT approaches into their application areas. Overall, 37/45 studies focused on typical SAT solving, 2 on QBFs, and 6 on 3SAT. Out of 45 articles, 9 focused on solving, 9 on finding MaxSAT, 7 on solver selection, 6 on feature extraction, 5 aimed to predict solubility and four contributions aimed at analyzing the performance of solvers. A small group of contributions aimed to do multiple things generated SAT problems and developed branching heuristics aimed at parameter selection or dependency analysis.
Figure 10(d) (see also Appendix B) takes a closer look at the topic of synthesis. 8/19 contributions aimed to synthesize a model or specification, 4/19 aimed to learn a loop invariant, 3/19 aimed to repair models, and 3/19 targeted general invariant learning. One contribution targets the generation of annotations for the verification of JML code.
(a) Quantity of targeted FM types
(b) Quantity of targeted theorem proving topics
(c) Quantity of targeted SAT topics
(d) Quantity of targeted synthesis topics
Figure 10: Quantity of FM approaches with a detailed view on the most relevant contributions types
Detailed view on theorem proving
A closer inspection is warranted as TP makes up the most significant portion of contributions. For the studies concerned with TP, we distinguished between higher-order logic (47/85) and first-order logic (38/85). We further classified them between their automation levels: fully automatic TP (74/85), interactive TP (10/85), and contributions employing both (1/85). This granularity gives rise to additional observations about the problem structure. Figure 11(a) (see also Appendix B) shows that the majority of contributions (26/85) were made in the area of premise selection for automatic theorem proving, followed by proof search (21/85) and proof synthesis (18/85).
Figure 11(b) (see also Appendix B) shows that of the 38 first-order entries, there is a tight focus on premise selection (20/38), proof search (8/38) and proof synthesis (7/38). Although not as drastic, a similar spread can be observed for the higher-order contributions. Premise selection has 7/47 entries, proof search 13/47, and proof synthesis 11/47 entries. 4/47 entries focused on tactics prediction.
(a) Division of TP applications into first-order logic (FOL) and higher-order logic (HOL).
(b) Automatization grade against the type of TP subcategory
Figure 11: Theorem proving in details
Intersecting RQ 2.2 with RQs 1.1 and 1.3
Figure 12(a) (see also Appendix B) shows the different types of FM over the years. We can see no particular concentration on one year, similar to Fig. 9, However, there may be some trends. SAT had two solid years in 2020 and 2022, while TP had a slow decline; contributions to model checking and synthesis increased. Nonetheless, given the small time frame of five years, Any actual trends might be invisible to us.
In Fig. 12(b) (see also Appendix B), we showcase the distribution of FM types over contribution types. While all FM techniques found AI applications in the last five years, we now see a fragmented distribution of specific contribution types per discipline. In general, for TP, the different types of available contributions are the broadest, while for model checking, we are restricted to methodologies and one tool. The notable absence of tool improvements outside of TP might indicate developments in the respective subfields. There might be no developments that deem an extension of tools necessary, or no relevant tools can be extended. However, the underlying reasons for this observation are out of the scope of this mapping study. For the synthesis of formal models, we see a slightly better situation regarding a broader landscape of contribution types. Overall, however, the absence of case studies and benchmarks deserves attention and might highlight that the field is still in the early stages of development.
(a) Distribution of FM techniques over the years
(b) Distribution of FM techniques over contribution types. Studies with multiple contribution types are considered for each distinct contribution type individually.
Figure 12: Quantity of FM approaches and comparison with time frame and contribution type
5.2.3 RQ 2.3: What is the distribution of AI types in the different FM application areas?
In Fig. 13, we can see several overlays of FM applications with FM techniques in different granularity. Figure 13(a) (see also Appendix B) is the most abstract. Here, we can see that the bulk of the theorem-proving techniques apply NNs (38), RL (16), and NLP (12). SAT solvers mainly utilize NNs (24). Model checking frequently uses EA (6) and NN (5). The area of synthesis is mainly divided between NNs (5), RL (4), and NLP (6) as well.
Taking a closer look at TP in Fig. 13(b) (see also Appendix B), we can see that NN is primarily used for the premise, axiom, and clause selection (18) proof search (7) and proof synthesis (10) RL techniques are often used for proof search as well (8). With three exceptions for heuristic selection, premise selection, or proof search, EA is absent from the area of theorem proving.
Focusing on SAT in Fig. 13(c) (see also Appendix B), we can see that NN is dominant in most subcategories, except for portfolio selection and branching heuristics. EA has higher relevance for MaxSat.
Finally, taking a closer look at the synthesis topic in Fig. 13(d), we can see that NLP is relevant for the model generation (see also Appendix B) and the generation of verification annotations. NN and RL are mainly used for (loop) invariant learning. For the repair of formal models, the contributions consider multiple techniques.
(a) FM techniques with applied AI techniques
(b) Theorem proving techniques with applied AI techniques
(c) SAT with applied AI techniques
(d) Synthesis with applied AI techniques
Figure 13: Quantity of FM approaches and comparison with time frame and contribution type. Studies that applied multiple AI techniques types are considered for each distinct AI technique individually. Underrepresented techniques are aggregated as Other with a number in parenthesis indicating how many sub-techniques were included.
5.2.4 RQ 2.4: Are the studies’ employed data sets publicly available?
Lastly, we investigated the found data sets. Here, multiple observations took place.
First, we searched for high-quality data sets, enabling other researchers to conduct their research efficiently. As pointed out earlier, we found 21 data sets. In Fig. 14 (see also Appendix B), we see the respective scopes of the data sets. Of the 21 data sets, 15 focused on TP, while 3 aimed at synthesis, 1 at SAT, 1 at SMT, and 1 at MC.
However, we were also interested in the data sets, which were not novel contributions but instead found reuse in our found studies. For this, we skimmed all our studies for data set mentions. As it turns out, of the 189 articles, 124 used external data sets. The 65 remaining studies generate random samples or do not mention their data source. Noteworthy is that when random data is used, the generated data is seldomly shared, so the reproduction of results is hindered.
The datasets found for the 124 contributions are largely heterogenic. This means that most studies utilized different datasets. However, we could find some data repositories that had multiple usages. The Mizar Library [^6] https://mizar.uwb.edu.pl/library/ (and associated libraries like, e.g., MPTP[47]) saw 23 usages, TPTP Problem Library [^7] https://tptp.org/TPTP/ has 7 mentions, The ProB machine library [^8] https://prob.hhu.de/w/index.php?title=Download got 5 mentions, HoL [^9] https://isabelle.in.tum.de/library/ 5, SMTLib [^10] https://smt-lib.org 5, SATComp [^11] https://satcompetition.github.io 5, SATLib [^12] https://www.cs.ubc.ca/~hoos/SATLIB/benchm.html 4, CoqGym [^13] https://github.com/princeton-vl/CoqGym 4. The full list is available in the repository linked in Section 1.
Figure 14: Overview of the distribution of data sets over FM domains
6 Discussion
In the following, we will evaluate the results and conclude to answer our research questions. Further, we will outline lessons learned to provide actionable insights for conducting future studies in the area.
6.1 RQ 1: Demographics of the research area
In RQs 1.1, 1.2 and 1.3, we investigated the publication timeline for possible trends, and assessed the qualitative nature of the studies, i.e., the maturity of the publication format as well as their level of postulated contribution, following our taxonomies as defined in Section 2.4. While we were able to analyze RQs 1.1 and 1.2 over the full corpus of 457 publications, we restricted ourselves to the processed corpus of 189 publications from the last five years for content assessment as outlined in Section 4.5. Consequently, RQ 1.3 is limited to these 189 publications.
For RQ 1.1, we see a general growth in the research field. However, from the presented data points it cannot be inferred alone whether this trend follows the overall trend of AI’s rise in popularity or if it is independent of it.
For RQ 1.2, we see a strong focus on conference papers, short papers, and workshops. Together with the singular book chapter, this indicates an emerging field, as there is a tendency to rely on publication methods supporting short review cycles. This is further supported by the fact that the only book chapter was a case study. Contrasting this with RQ 1.1, we assume that the research interest will at least not plummet in the foreseeable future but might plateau for the time being.
For RQ 1.3, we can see that the bulk of works lay within the area of methodologies. The reason for this could be the current paper requirements, requesting a novel approach and some evaluation within the paper. Another reason might be AI’s applicative nature, which often allows it to serve as a support system to automate parts of the work. We also want to highlight the possibility that we primarily introduced methodology papers based on the nature of our search query, which focuses on the application of AI in the FM domain. However, due to rigorous snowballing, we are confident we correctly assessed the state of research in this field.
What is more interesting is what we do not see: many case studies and benchmarks. Recalling Fig. 6, we only see two publications providing case studies and two providing benchmarks. This may hint at a further gap in maturity, especially as these contribution types are primarily present in journal and book contributions. One would assume more such publications if a set of robust solutions for common problems were readily available so that researchers could compare their effectiveness against each other. It also highlights the need for proper and uniform benchmark sets within the community which can be shared between different research groups.
6.2 RQ 2: Content type of contributions
In RQ 2, we were interested in how AI application is distributed in FM. For this, we posed RQs 2.1, 2.4, 2.2 and 2.3 to investigate which AI techniques find appreciation in the community, which FM areas already are subject to AI application, if certain AI techniques are predominantly used in specific FM fields, and whether there exist publicly available data sets which can be used in further research.
For RQ 2.1, we saw in Fig. 7 Most studies use NNs, followed by RL, and NLP and EAs are far behind. From a paradigmatic point of view, it makes sense to see these four on top, as they all attempt to solve a different problem setting. NNs are a stand-in for classification and regression tasks, RL is applied for learning behavioral policies, NLP is processing texts resembling natural language or code by extension, and EAs excel in optimization problems. However, it is interesting to see NNs being so far ahead of the rest with 83 publications utilizing them. Especially, since they are not the only algorithms suitable for classification and regression tasks, but competing algorithms such as KNN, SVM, or decision trees are, at least as solely applied AI techniques, vastly underrepresented. These algorithms, which we categorized as classical approaches in Section 2.3, are, however, not forgotten, as we can see in publications which contrast multiple different AI algorithms (Fig. 7(b)). This focus on NNs when only one technique is used can have multiple underlying reasons. First, implementation, training, and fine-tuning of NNs is a time-consuming task, and might take up the resources needed to investigate further algorithms. Second, we have seen in Fig. 3 that the research area gained traction in the early 2010s. As this coincides with the advent of deep learning [48], this might suggest the need for NNs to overcome the complexities the area of FM offers, and that more classical ML algorithms are not suitable for most tasks. Third, there might exist a bias towards using NNs from the get-go in the community instead of applying more straightforward techniques first.
However, the prominence of more classical algorithms in the case of multiple AI techniques showcases that the community seems to be aware that other algorithms might be suitable for their problems as well. This is especially amplified by Fig. 8 showing that most frequently, it is methodologies that rely on various techniques to solve problems.
Regarding maturity, Fig. 9 indicates no noteworthy increase in techniques besides NLP. For 2023, we established insufficient data points, so we may see a rise in NLP-based contributions, which would fit a more general advent of NLP-related contributions due to the rise of transformer architectures and LLMs.
For RQ 2.2, we have seen in Fig. 10(a) that TP and SAT are the predominant FM areas in which AI is applied, making up 68.8 % of all publications in our five-year corpus. We explain this by the fact that both areas are also of interest to non-FM communities. A detailed view of the authors’ origin domains would be necessary to verify this assumption. However, this is still beneficial for the FM community as a whole. For instance, Fig. 11(a) shows a strong focus on FM-relevant daily use cases. Figure 11(b) reveals much current work regarding first-order logic on the especially relevant FM topics of premise selection. While this is a good trend, complex FM models rely on complicated data types and require sophisticated reasoning.
In regards to SAT, we can take away from Fig. 10(c) that much focus lies on SAT solving directly, while other authors also attempt to predict satisfiability. The latter seems to stand in contradiction with the rigorousness needed in FM, as the probabilistic nature of AI approaches lacks the certainty of the predicted satisfiability as a whole. This can be a product of cross-fertilization, as SAT is also relevant for non-FM communities.
Figure 10(d) shows that the interest in synthesis is low compared to other topics and primarily focuses on model generation. Given the recent rise of generative AI approaches, we see room for promising research in this area in the future.
A similar low interest seems to be in model checking. Figure 12(b) shows a relatively low number of contributions in that area. The direct comparison with the TP field is interesting, as MC also contributes very little in tool and tool contributions, hinting at less application-focused research. Figure 12(a) shows a relatively consistent interest in recent years, while Fig. 12(b) emphasizes how little is done for current practitioners in terms of tool (improvements), benchmarks, training data, and case studies.
For RQ 2.3, we saw in Fig. 13(a) that NNS is the undeniable favorite in almost all FM areas investigated. We already discussed why NNs might be more popular than other algorithms suitable for the same problem tasks above for RQ 2.1. However, this suggests again how the FM community mainly focuses on classification or regression problems to solve rather than execution policies (RL) or direct processing of texts or code (NLP). The slight exception we see is in the area of model checking, which focuses most on EAs, suggesting a more substantial prevalence of optimization problems in this area. From building the larger corpus of 457 studies, we know, however, from sighting various titles that we missed many EA papers due to our 5-year restriction. Presumably, EA was more prominently used before deep learning took off.
An interesting observation is the substantial absence of any data mining-related results, which might solve other problems still not covered by NNs, RL, NLP, or EA. In the 189 investigated studies, only three applied data mining techniques. It is again worth noting that we only investigated the years 2019–2023. Hence, data mining may have been more widespread before that, much like EA. More thorough data mining would be needed for a deeper insight, and only a secondary study would be required.
Finally, for RQ 2.4, we found 19 data set (or training environment) contributions in our corpus of 189 studies. Considering the two separate, data set-only publications which did not pass IC 1 (see Section 4.5), this leaves us with 21 total data set contributions.
Together with the analyzed high heterogeneity of the used datasets, as already mentioned in Section 5.1.3, there may be a general underlying issue. There seems to be no defined benchmark against which to run new insights. While datasets are publicly available for TP and SAT solving, the number of usable, established, and available benchmark data sets is minuscule for other areas.
Generally, we see a current trend where researchers use datasets as they see fit or even generate data without referencing the procedure or providing the data set afterward. This trend fits the overall trend of AI being only weakly reproducible [49]. However, this trend endangers the whole research branch, as reproducibility is a core part of science.
Concluding, we see a strong need for unified defaults to not only allow reproducibility but also to introduce familiar benchmarks into the field. Possible issues hindering such a development might be found in the various existing formalisms, which are not necessarily used or even supported by individual research groups. Such boundaries, of course, weaken the impact a data set publication might have onto researchers working with differing formalisms.
6.3 Suggested research directions
From the given analysis, we can derive five areas of growth potential, which we outline below.
Development of unified benchmarks and data sets
As pointed out in our discussion to RQ 2.4, we are concerned with the lack of established data sets and benchmarks in the FM su-communities. While for TP, a good amount of contributions exist already tackling this problem, the other areas seemingly fall short. This creates room for reproducibility issues between research groups and hinders a more precise comparison between results as established benchmarks could provide.
Developing strong baselines might not only increase comparability between different methods but might also reduce the entry hurdle for new researchers due to publicly available data sets. The main challenge is to create data sets which can be used for different formalisms. This would allow smaller communities to still benefit from available data which might be posed in a more popular formalism while furthering research on their area.
An orientation on how such a unified benchmark environment could look like would be OpenAIGym [50], a Toolkit for RL to foster research by enhancing reproducibility.
Investigate the benefits of data mining for FM
In the age of big data, we were expecting more studies that apply data mining techniques, be it for feature engineering or performance analytical approaches. This was not the case. While it might very well be that the kind of problems solvable by data mining are not relevant to the FM communities, it is also possible that potential benefits are not well-known.
We propose to research if and how data mining approaches can benefit FM. First, this would highlight potential benefits to the community. Second, due to the investigative nature of data mining, this might strengthen our insides in our tools and best practices. Consequently, this might further the quality of other AI applications in FM due to new knowledge.
Application of LLMs and generative AI
While LLMs, at this point, are established tools that found their way into daily applications, we only found little use of them so far in our corpus of studies. This can be due to their relatively high employment costs. Another reason could be that respective studies are still in the process of being published and were hence invisible to our systematic mapping study. Nevertheless, we see strong potential for generative AI approaches, especially in the areas of synthesis and auto-formalization. However, also simple test-case generation or the generation of documentation for existing models seem promising starting points to build trust in the process.
We envision substantial benefits from applying generative AI during FM development and point out the need for more research. One can easily envision tool chains of automatic generation, checking, and documentation of formal models with reproducible and provable results. However, we also see the need for more exploration regarding how generative AI can be helpful besides text production. For instance, how could generative AI be leveraged in proof mining? What stages in the formal development process can benefit from generated inputs, and what types could and should be generated? Pure generation of code and text might not be the ceiling for FM.
Increase the potentials of AI in model checking
Another underrepresented topic is model checking. Anecdotally, we know model checking had much EA applied in the past while it was a seemingly unpopular application area for AI in our five-year corpus. However, we see a significant potential to revisit respective research with a modern lens. Especially in the context of bounded model checking (BMC) or statistical model checking (SMC) and faulty state prediction. BMC cuts off parts of a model’s state space for performance and computability. SMC checks states until a certain confidence in the absence of faulty states is reached. Both approaches already do not fully traverse the state space and can only indicate the absence of faults.
AI-enforced model checking could alleviate BMC, SMC, and model checking as a whole by predicting potential faulty states or needed search depths, learning search policies via RL to reach faulty states in a more targeted manner, or predict equivalence classes of states for search space reduction.
7 Threats to validity
A study of this type and extent can only be conducted with the possibility of encountering threats that threaten validity. We orient ourselves at Zhou et al. Zhou et al. [30] for addressing the most critical ones.
Internal validity
A potential threat to the conclusion cloud is our selection of articles and our data source. To ensure the quality of our data, we carefully drafted our IC/EC and in cases where we found them not yet sufficient, we enhanced them. Two experienced researchers carefully read the publications.
Our data source was four famously known databases containing the relevant literature [51]. Due to the extensive snowballing, we made sure that the potential for blind spots, i.e., publications that are only listed by one engine or none at all, is minimized. This also holds for the case of WoS, where different result sizes are obtained. The divergence between both result sets was marginal, so we are confident excessive snowballing alleviates it. The snowballing also alleviated some of our rigorousness by applying IC/EC very early on in our search process. While we reduced our initial tally of 1492 publications to only 89, a reduction of two orders of magnitude, snowballing helped us to ensure recapturing significant contributions into our database again.
Construct validity
Another threat is the use of search terms. We used PICO as Kitchenham et al. [10] suggest to select the appropriate search term. Nevertheless, we know it is impossible to cover all types of FM, AI, and all their acronyms. Therefore, the aim was to conduct an exhausting snowballing procedure to include all relevant publications.
The amount of snowballing results was unsurprisingly large, as snowballing does not suffer from a compromise for a search query. This is a limitation already pointed out by Wohlin [46].
Conclusion validity
Grey literature was not considered for a variety of reasons. First, we did not encounter large quantities of grey literature during our snowballing. Thus, We can represent current research directions even without considering them. Second, We were concerned about the feasibility of the provided workshop and short papers. As we only did a soft skimming of the content, we can not offer a quality and feasibility assessment of the included contributions. As we aim to provide the reader with only high-quality studies, we decided to require official publications.
8 Related work
While the lack of available literature partially motivated this work, there are noteworthy related works. One of the first contributions one finds when investigating the topic is the one of Amrani et al. [52]. Here, the authors came to a similar conclusion as we did about the complicated nature of the investigated landscape of available literature. Therefore, they also resorted to a complex search string. The evaluation and focus, however, are different. The authors limit themselves to an excerpt of initial results and only conduct backward snowballing, leaving room for whether the corpus is complete. Additionally, the research questions are more specifically tailored to individual FM topics, while our work aims to give a general overview.
Wang et al. [53] aims to give a general taxonomy of learning-based FM. One core contribution is that they divide the topic into two general directions: the learning of formal specifications and learning for formal verification. They proceeded to structure the literature found by individuals underneath this taxonomy. Compared to their work, our work aims to give a general introduction to the cross-section of the two fields while the authors dive into the specifics of learning algorithms for formal verification.
Solvers and solving
Large quantities of work exist to evaluate the use of AI to enable solvers. A general overview was given by Ganesh et al. [54]. In the context of SAT, multiple works exist. A general overview with comments on the use of AI was presented by Kilani et al. [55]. Explicit SAT solving was subject to two studies [56, 57]. Configuration and benchmarking of SAT solvers were investigated in three contributions [58, 59, 60].
During our search, we found many contributions that targeted the general constraint satisfaction problem. However, we excluded these results as they are only loosely connected with FM. Amadini et al. [61] conducted a general evaluation of portfolio approaches for CSP problems. Popescu et al. [62] gives an overview of machine learning techniques for CSP problems. While Yang et al. [63] conducts a detailed survey of intelligent optimization algorithms for satisfaction problems.
Proving
Multiple works aim to overview the topic within the math and automated reasoning community. Urban and Vyskočil [64] give an overview of AI-supported automated theorem proving. Blanchette et al. [65] surveyed ML-supported axiom selection, and two years later, Blanchette et al. [66] surveyed the rising interest in automatizing reasoning over proofs. England [67] published a survey on ML for symbolic reasoning. Tran et al. [68] gave a short overview of the emerging field of NLP for premise selection.
Program analysis
For general program analysis, there are works for runtime prediction by Hutter et al. [69]. Kumazawa et al. [70] published a survey on applying swarm intelligence (a flavor of EAs) for software verification. Bennaceur and Meinke [71] published on which ML techniques may be suitable for software analysis.
Model/specification generation
Brunello et al. [72] and Buzhinsky [73] published surveys on creating temporal logic from natural language, while Fuggitti and Chakraborti [74]. Szegedy [75] discusses the topic of auto formalization of natural language requirements.
Others
Less prominent were the topics of model repair, where Barriga et al. [76] published a longer article about the state of the art. Haltermann and Wehrheim [77], gives an overview of invariant generation. Pan and Mishra [78] surveyed hardware vulnerability analysis via ML, a field to which we found no contributions. For model checking, Besbas et al. [79] gave a brief literature review of four pages.
9 Conclusion
In this systematic mapping study, we present the results of our work on assessing the quantity of research in applying artificial intelligence to formal methods. Overall, we found 457, from which we selected 189 for a closer investigation. Concluding from this investigation, we see the current trend is yet to mature, as many contributions are making some practical application. However, only a few studies aim to create theoretical groundwork, benchmarks, or case studies. Furthermore, we see a focus on theorem proving and SAT solving. Model checking and model synthesis are underrepresented compared to this. Later may be surprising with the advantage of LLM that becomes precise in creating output after a user’s desire. Most work uses some neural nets, reinforcement learning, or a combination of multiple approaches.
Appendix A Tables for RQ 1
Entries grouped by years
Year Articles Count
[…]
Year Articles Count
[…]
2020 [112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159] 48
2021 [160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198] 39
2022 [199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249] 51
2023 [250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 60, 263, 264, 265, 266, 267, 268, 269] 21
Entries grouped by type of publication venue
Entry type Articles Count
[…]
Entry type Articles Count
[…]
Book chapter [233] 1
Workshop/short paper [94, 95, 96, 99, 105, 120, 132, 137, 138, 148, 154, 156, 160, 162, 166, 173, 175, 176, 192, 200, 214, 223, 228, 236, 238, 246, 259, 264, 268] 29
Conference paper [81, 83, 84, 85, 88, 89, 90, 91, 92, 93, 97, 101, 102, 103, 104, 106, 107, 113, 115, 116, 117, 118, 119, 121, 123, 124, 125, 126, 127, 128, 129, 130, 134, 139, 140, 141, 142, 144, 145, 146, 149, 150, 151, 152, 153, 155, 157, 158, 161, 163, 164, 167, 168, 169, 171, 172, 174, 177, 178, 179, 181, 182, 183, 184, 185, 186, 189, 190, 191, 193, 194, 195, 196, 197, 198, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 216, 217, 218, 220, 221, 222, 224, 225, 227, 229, 230, 231, 232, 234, 235, 239, 240, 242, 243, 244, 245, 247, 252, 254, 256, 260, 261, 263, 265, 266, 267, 269, 110, 111, 159] 122
Entries grouped by type of contribution
Contribution type Articles Count
[…]
Contribution type Articles Count
[…]
Tool improvement [84, 85, 91, 144, 178, 179, 189, 208, 224, 254] 10
Idea [242] 1
Approach [107, 116] 2
Framework [167, 193, 203, 217, 219, 247] 6
Methodology [81, 82, 86, 87, 88, 93, 94, 95, 96, 98, 99, 101, 103, 105, 106, 108, 109, 113, 114, 117, 118, 119, 121, 123, 124, 127, 128, 129, 133, 134, 135, 136, 137, 140, 141, 142, 143, 147, 148, 149, 151, 154, 155, 156, 157, 158, 160, 161, 162, 163, 165, 166, 168, 169, 170, 173, 175, 176, 177, 181, 182, 184, 185, 186, 187, 188, 194, 196, 197, 198, 199, 200, 201, 202, 204, 205, 206, 207, 212, 213, 214, 215, 216, 218, 220, 221, 223, 225, 226, 227, 229, 231, 234, 235, 236, 237, 238, 239, 240, 243, 245, 248, 249, 250, 251, 252, 255, 257, 259, 260, 261, 262, 60, 264, 266, 268, 269, 111] 118
Case study [233, 253] 2
Benchmark [112, 150] 2
Training data [92, 145, 146, 241, 246] 5
Mult. contributions [80, 89, 90, 97, 100, 104, 125, 126, 130, 132, 138, 139, 171, 172, 174, 180, 183, 195, 209, 211, 222, 228, 232, 244, 256, 258, 267, 110, 159] 29
Entries grouped by type of contribution and publication venue
Contribution type Entry type Articles Count
[…]
Contribution type Entry type Articles Count
[…]
Workshop/short paper [120, 192] 2
Conference paper [83, 102, 115, 152, 164, 190, 191, 210, 230, 263] 10
Tool improvement Conference paper [84, 85, 91, 144, 178, 179, 189, 208, 224, 254] 10
Idea Conference paper [242] 1
Approach Conference paper [107, 116] 2
Framework Journal paper [219] 1
Conference paper [167, 193, 203, 217, 247] 5
Methodology Journal paper [82, 86, 87, 98, 108, 109, 114, 133, 135, 136, 143, 147, 165, 170, 187, 188, 199, 213, 215, 226, 237, 248, 249, 250, 251, 255, 257, 262, 60] 29
Workshop/short paper [94, 95, 96, 99, 105, 137, 148, 154, 156, 160, 162, 166, 173, 175, 176, 200, 214, 223, 236, 238, 259, 264, 268] 23
Conference paper [81, 88, 93, 101, 103, 106, 113, 117, 118, 119, 121, 123, 124, 127, 128, 129, 134, 140, 141, 142, 149, 151, 155, 157, 158, 161, 163, 168, 169, 177, 181, 182, 184, 185, 186, 194, 196, 197, 198, 201, 202, 204, 205, 206, 207, 212, 216, 218, 220, 221, 225, 227, 229, 231, 234, 235, 239, 240, 243, 245, 252, 260, 261, 266, 269, 111] 66
Case study Journal paper [253] 1
Book chapter [233] 1
Benchmark Journal paper [112] 1
Conference paper [150] 1
Training data Journal paper [241] 1
Workshop/short paper [246] 1
Conference paper [92, 145, 146] 3
Mult. contributions Journal paper [80, 100, 180, 258] 4
Workshop/short paper [132, 138, 228] 3
Conference paper [89, 90, 97, 104, 125, 126, 130, 139, 171, 172, 174, 183, 195, 209, 211, 222, 232, 244, 256, 267, 110, 159] 22
Appendix B Tables for RQ 2
Entries grouped by type of AI type used
AI technique Articles Count
[…]
AI technique Articles Count
[…]
RF [117, 201] 2
SVM [133] 1
KNN [122, 125] 2
NN [83, 90, 92, 96, 97, 99, 102, 104, 106, 120, 121, 123, 124, 126, 129, 134, 135, 137, 139, 140, 141, 146, 150, 151, 152, 155, 158, 163, 165, 169, 170, 174, 175, 177, 178, 179, 182, 190, 192, 193, 195, 196, 203, 205, 206, 209, 210, 211, 218, 219, 220, 222, 227, 228, 231, 232, 233, 234, 237, 239, 242, 254, 257, 258, 259, 262, 263, 268, 110, 159] 70
RL [94, 98, 100, 101, 115, 118, 127, 130, 132, 142, 148, 149, 161, 167, 168, 171, 176, 184, 191, 199, 207, 224, 230, 235, 240, 241, 250, 253, 256, 260, 264, 267] 32
EA [81, 107, 108, 109, 147, 188, 194, 198, 200, 204, 213, 243, 248, 269] 14
NLP [116, 131, 145, 160, 162, 183, 186, 197, 202, 217, 223, 229, 236, 238, 245, 246] 16
Automaton learning [138, 249] 2
Bayesian inference [113, 136, 221] 3
Clustering [86, 114, 244] 3
Data mining [88, 143] 2
Mult. AI techniques [80, 82, 85, 87, 89, 93, 105, 112, 119, 128, 156, 164, 166, 172, 173, 180, 187, 189, 208, 212, 215, 226, 251, 252, 261, 266, 111] 27
Custom [103, 154, 157, 185, 214, 247, 255, 60] 8
Entries that use multiple AI types
AI technique Articles Count
[…]
AI technique Articles Count
[…]
SVM [85, 93, 166, 180, 215, 226] 6
KNN [85, 87, 166, 172, 226] 5
NN [80, 89, 112, 128, 173, 180, 187, 189, 208, 212, 251, 252, 111] 13
EA [187] 1
Clustering [105, 156, 212] 3
Data mining [119] 1
Logistic models [80, 87, 93, 112, 215, 226] 6
LR [85, 87, 166] 3
[82] 1
Ridge regression [166, 226, 251] 3
Naive bayes [93, 112, 215, 226] 4
Trees [80, 82, 85, 105, 112, 156, 164, 166, 172, 189, 226, 111] 12
Nearest centroid [226] 1
Passive aggressive [226] 1
Stochastic gradient [226] 1
LLM [261, 266] 2
Bagging [85] 1
Adaboost [85, 164] 2
Gaussian process [226] 1
MCTS [261] 1
SMAC [105] 1
Entries grouped by AI types and contribution type
Contribution type AI technique Articles Count
[…]
Contribution type AI technique Articles Count
[…]
NN [83, 102, 120, 152, 190, 192, 210, 263] 8
RL [115, 191, 230] 3
NLP [131] 1
Mult. AI techniques [164] 1
Tool improvement DT [84, 91, 144] 3
NN [178, 179, 254] 3
RL [224] 1
Mult. AI techniques [85, 189, 208] 3
Idea NN [242] 1
Approach EA [107] 1
NLP [116] 1
Framework NN [193, 203, 219] 3
RL [167] 1
NLP [217] 1
Custom [247] 1
Methodology DT [95, 181, 216, 225] 4
RF [117, 201] 2
SVM [133] 1
NN [96, 99, 106, 121, 123, 124, 129, 134, 135, 137, 140, 141, 151, 155, 158, 163, 165, 169, 170, 175, 177, 182, 196, 205, 206, 218, 220, 227, 231, 234, 237, 239, 257, 259, 262, 268] 36
RL [94, 98, 101, 118, 127, 142, 148, 149, 161, 168, 176, 184, 199, 207, 235, 240, 250, 260, 264] 19
EA [81, 108, 109, 147, 188, 194, 198, 200, 204, 213, 243, 248, 269] 13
NLP [160, 162, 186, 197, 202, 223, 229, 236, 238, 245] 10
Automaton learning [249] 1
Bayesian inference [113, 136, 221] 3
Clustering [86, 114] 2
Data mining [88, 143] 2
Mult. AI techniques [82, 87, 93, 105, 119, 128, 156, 166, 173, 187, 212, 215, 226, 251, 252, 261, 266, 111] 18
Custom [103, 154, 157, 185, 214, 255, 60] 7
Case study NN [233] 1
RL [253] 1
Benchmark NN [150] 1
Mult. AI techniques [112] 1
Training data NN [92, 146] 2
RL [241] 1
NLP [145, 246] 2
Mult. contributions KNN [125] 1
NN [90, 97, 104, 126, 139, 174, 195, 209, 211, 222, 228, 232, 258, 110, 159] 15
RL [100, 130, 132, 171, 256, 267] 6
NLP [183] 1
Automaton learning [138] 1
Clustering [244] 1
Mult. AI techniques [80, 89, 172, 180] 4
Entries grouped by AI types and publication years
AI technique Year Articles Count
[…]
AI technique Year Articles Count
[…]
2020 [144] 1
2021 [181] 1
2022 [216, 225] 2
RF 2020 [117] 1
2022 [201] 1
SVM 2020 [133] 1
KNN 2020 [122, 125] 2
NN 2019 [83, 90, 92, 96, 97, 99, 102, 104, 106, 110] 10
2020 [120, 121, 123, 124, 126, 129, 134, 135, 137, 139, 140, 141, 146, 150, 151, 152, 155, 158, 159] 19
2021 [163, 165, 169, 170, 174, 175, 177, 178, 179, 182, 190, 192, 193, 195, 196] 15
2022 [203, 205, 206, 209, 210, 211, 218, 219, 220, 222, 227, 228, 231, 232, 233, 234, 237, 239, 242] 19
2023 [254, 257, 258, 259, 262, 263, 268] 7
RL 2019 [94, 98, 100, 101] 4
2020 [115, 118, 127, 130, 132, 142, 148, 149] 8
2021 [161, 167, 168, 171, 176, 184, 191] 7
2022 [199, 207, 224, 230, 235, 240, 241] 7
2023 [250, 253, 256, 260, 264, 267] 6
EA 2019 [81, 107, 108, 109] 4
2020 [147] 1
2021 [188, 194, 198] 3
2022 [200, 204, 213, 243, 248] 5
2023 [269] 1
NLP 2020 [116, 131, 145] 3
2021 [160, 162, 183, 186, 197] 5
2022 [202, 217, 223, 229, 236, 238, 245, 246] 8
Automaton learning 2020 [138] 1
2022 [249] 1
Bayesian inference 2020 [113, 136] 2
2022 [221] 1
Clustering 2019 [86] 1
2020 [114] 1
2022 [244] 1
Data mining 2019 [88] 1
2020 [143] 1
Mult. AI techniques 2019 [80, 82, 85, 87, 89, 93, 105, 111] 8
2020 [112, 119, 128, 156] 4
2021 [164, 166, 172, 173, 180, 187, 189] 7
2022 [208, 212, 215, 226] 4
2023 [251, 252, 261, 266] 4
Custom 2019 [103] 1
2020 [154, 157] 2
2021 [185] 1
2022 [214, 247] 2
2023 [255, 60] 2
Entries grouped by FM types
FM technique Articles Count
[…]
FM technique Articles Count
[…]
SMT [89, 95, 119, 161, 163, 164, 171, 185, 186, 216, 251, 263, 264] 13
TP [83, 84, 87, 88, 91, 92, 94, 96, 97, 99, 102, 105, 106, 107, 115, 120, 122, 123, 124, 125, 126, 129, 130, 131, 132, 136, 138, 139, 143, 144, 145, 146, 148, 149, 152, 154, 155, 156, 165, 167, 168, 172, 173, 174, 175, 176, 177, 178, 179, 180, 183, 184, 187, 188, 189, 190, 191, 196, 197, 207, 208, 209, 214, 217, 218, 221, 223, 224, 225, 228, 229, 232, 236, 237, 238, 241, 245, 250, 253, 254, 257, 258, 261, 266, 111] 85
MC [81, 82, 108, 109, 113, 118, 166, 195, 198, 205, 220, 226, 230, 248, 249, 252, 268, 269] 18
Synthesis [80, 93, 116, 157, 160, 162, 192, 193, 199, 202, 210, 211, 215, 244, 246, 247, 256, 260, 267] 19
Other [100, 134, 135, 140, 182, 227, 231, 233, 243] 9
Entries grouped by TP application types
Application goal Articles Count
[…]
Application goal Articles Count
[…]
Premise selection [84, 91, 96, 124, 139, 154, 155, 156, 168, 173, 174, 177, 178, 188, 189, 196, 197, 209, 214, 217, 218, 221, 223, 237, 241, 254, 111] 27
Proof search [83, 92, 94, 115, 120, 122, 123, 125, 130, 136, 144, 145, 146, 165, 176, 187, 191, 224, 225, 250, 261] 21
Proof mining [143, 183] 2
Proof synthesis [97, 99, 126, 131, 138, 149, 175, 190, 207, 208, 228, 229, 232, 236, 238, 253, 257, 258] 18
Symbolic [184] 1
Formula synthesis [148] 1
Formula classification [179, 180] 2
Tactic prediction [88, 106, 132, 172] 4
Lemma name generation [152] 1
Portfolio solving [87] 1
Mult. Techniques [102, 129, 167, 245, 266] 5
Entries grouped by synthesis application types
Application goal Articles Count
[…]
Application goal Articles Count
[…]
Loop invariant learning [157, 192, 244, 267] 4
Model repair [80, 93, 215] 3
Models/Specification [116, 160, 199, 202, 210, 211, 246, 247] 8
Annotations [162] 1
Entries grouped by grade of TP automatization and TP application types
Application goal Automation grade Articles Count
[…]
Application goal Automation grade Articles Count
[…]
Premise selection Interactive [91] 1
Automatic [84, 96, 124, 139, 154, 155, 156, 168, 173, 174, 177, 178, 188, 189, 196, 197, 209, 214, 217, 218, 221, 223, 237, 241, 254, 111] 26
Proof search Automatic [83, 92, 94, 115, 120, 122, 123, 125, 130, 136, 144, 145, 146, 165, 176, 187, 191, 224, 225, 250, 261] 21
Proof mining Interactive [143] 1
Automatic [183] 1
Proof synthesis Automatic [97, 99, 126, 131, 138, 149, 175, 190, 207, 208, 228, 229, 232, 236, 238, 253, 257, 258] 18
Symbolic Automatic [184] 1
Formula synthesis Automatic [148] 1
Formula classification Interactive [179] 1
Automatic [180] 1
Tactic prediction Interactive [88, 132, 172] 3
Automatic [106] 1
Lemma name generation Interactive [152] 1
Portfolio solving Automatic [87] 1
Mult. Techniques Interactive [102, 167, 266] 3
Automatic [129] 1
Mixed [245] 1
Entries grouped by level of logic tackled by TP and TP application types
Application goal Formalism Articles Count
[…]
Application goal Formalism Articles Count
[…]
HOL [107] 1
Premise selection FOL [84, 91, 96, 124, 154, 156, 168, 173, 174, 177, 178, 189, 196, 209, 214, 221, 237, 241, 254, 111] 20
HOL [139, 155, 188, 197, 217, 218, 223] 7
Proof search FOL [83, 115, 136, 144, 145, 146, 165, 250] 8
HOL [92, 94, 120, 122, 123, 125, 130, 176, 187, 191, 224, 225, 261] 13
Proof mining HOL [143, 183] 2
Proof synthesis FOL [99, 190, 207, 208, 229, 236, 238] 7
HOL [97, 126, 131, 138, 149, 175, 228, 232, 253, 257, 258] 11
Symbolic HOL [184] 1
Formula synthesis HOL [148] 1
Formula classification FOL [179, 180] 2
Tactic prediction HOL [88, 106, 132, 172] 4
Lemma name generation HOL [152] 1
Portfolio solving HOL [87] 1
Mult. Techniques HOL [102, 129, 167, 245, 266] 5
Entries grouped by FM techniques and years
FM technique Year Articles Count
[…]
FM technique Year Articles Count
[…]
2020 [112, 114, 117, 121, 127, 128, 133, 137, 141, 142, 147, 150, 151, 158, 159] 15
2021 [169, 170, 181, 194] 4
2022 [200, 201, 203, 204, 206, 212, 213, 219, 222, 234, 235, 239, 240, 242] 14
2023 [255, 259, 262, 60] 4
SMT 2019 [89, 95] 2
2020 [119] 1
2021 [161, 163, 164, 171, 185, 186] 6
2022 [216] 1
2023 [251, 263, 264] 3
TP 2019 [83, 84, 87, 88, 91, 92, 94, 96, 97, 99, 102, 105, 106, 107, 111] 15
2020 [115, 120, 122, 123, 124, 125, 126, 129, 130, 131, 132, 136, 138, 139, 143, 144, 145, 146, 148, 149, 152, 154, 155, 156] 24
2021 [165, 167, 168, 172, 173, 174, 175, 176, 177, 178, 179, 180, 183, 184, 187, 188, 189, 190, 191, 196, 197] 21
2022 [207, 208, 209, 214, 217, 218, 221, 223, 224, 225, 228, 229, 232, 236, 237, 238, 241, 245] 18
2023 [250, 253, 254, 257, 258, 261, 266] 7
MC 2019 [81, 82, 108, 109] 4
2020 [113, 118] 2
2021 [166, 195, 198] 3
2022 [205, 220, 226, 230, 248, 249] 6
2023 [252, 268, 269] 3
Synthesis 2019 [80, 93] 2
2020 [116, 157] 2
2021 [160, 162, 192, 193] 4
2022 [199, 202, 210, 211, 215, 244, 246, 247] 8
2023 [256, 260, 267] 3
Other 2019 [100] 1
2020 [134, 135, 140] 3
2021 [182] 1
2022 [227, 231, 233, 243] 4
Entries grouped by FM techniques and contributions types
FM technique Entry type Articles Count
[…]
FM technique Entry type Articles Count
[…]
Workshop/short paper [137, 200, 259] 3
Conference paper [85, 90, 101, 103, 104, 117, 121, 127, 128, 141, 142, 150, 151, 158, 169, 181, 194, 201, 203, 204, 206, 212, 222, 234, 235, 239, 240, 242, 110, 159] 30
SMT Journal paper [251] 1
Workshop/short paper [95, 264] 2
Conference paper [89, 119, 161, 163, 164, 171, 185, 186, 216, 263] 10
TP Journal paper [87, 122, 131, 136, 143, 165, 180, 187, 188, 237, 241, 250, 253, 257, 258] 15
Workshop/short paper [94, 96, 99, 105, 120, 132, 138, 148, 154, 156, 173, 175, 176, 214, 223, 228, 236, 238] 18
Conference paper [83, 84, 88, 91, 92, 97, 102, 106, 107, 115, 123, 124, 125, 126, 129, 130, 139, 144, 145, 146, 149, 152, 155, 167, 168, 172, 174, 177, 178, 179, 183, 184, 189, 190, 191, 196, 197, 207, 208, 209, 217, 218, 221, 224, 225, 229, 232, 245, 254, 261, 266, 111] 52
MC Journal paper [82, 108, 109, 226, 248, 249] 6
Workshop/short paper [166, 268] 2
Conference paper [81, 113, 118, 195, 198, 205, 220, 230, 252, 269] 10
Synthesis Journal paper [80, 199, 215] 3
Workshop/short paper [160, 162, 192, 246] 4
Conference paper [93, 116, 157, 193, 202, 210, 211, 244, 247, 256, 260, 267] 12
Other Journal paper [100, 135] 2
Book chapter [233] 1
Conference paper [134, 140, 182, 227, 231, 243] 6
Entries grouped by FM and AI techniques
FM technique AI technique Articles Count
[…]
FM technique AI technique Articles Count
[…]
RF [117, 201] 2
SVM [133] 1
NN [90, 104, 121, 137, 141, 150, 151, 158, 169, 170, 203, 206, 219, 222, 234, 239, 242, 259, 262, 110, 159] 21
RL [98, 101, 127, 142, 235, 240] 6
EA [147, 194, 200, 204, 213] 5
Clustering [86, 114] 2
Mult. AI techniques [85, 112, 128, 212] 4
Custom [103, 255, 60] 3
SMT DT [95, 216] 2
NN [163, 263] 2
RL [161, 171, 264] 3
NLP [186] 1
Mult. AI techniques [89, 119, 164, 251] 4
Custom [185] 1
TP DT [84, 91, 144, 225] 4
KNN [122, 125] 2
NN [83, 92, 96, 97, 99, 102, 106, 120, 123, 124, 126, 129, 139, 146, 152, 155, 165, 174, 175, 177, 178, 179, 190, 196, 209, 218, 228, 232, 237, 254, 257, 258] 32
RL [94, 115, 130, 132, 148, 149, 167, 168, 176, 184, 191, 207, 224, 241, 250, 253] 16
EA [107, 188] 2
NLP [131, 145, 183, 197, 217, 223, 229, 236, 238, 245] 10
Automaton learning [138] 1
Bayesian inference [136, 221] 2
Data mining [88, 143] 2
Mult. AI techniques [87, 105, 156, 172, 173, 180, 187, 189, 208, 261, 266, 111] 12
Custom [154, 214] 2
MC NN [195, 205, 220, 268] 4
RL [118, 230] 2
EA [81, 108, 109, 198, 248, 269] 6
Automaton learning [249] 1
Bayesian inference [113] 1
Mult. AI techniques [82, 166, 226, 252] 4
Synthesis NN [192, 193, 210, 211] 4
RL [199, 256, 260, 267] 4
NLP [116, 160, 162, 202, 246] 5
Clustering [244] 1
Mult. AI techniques [80, 93, 215] 3
Custom [157, 247] 2
Other NN [134, 135, 140, 182, 227, 231, 233] 7
RL [100] 1
EA [243] 1
Entries grouped by TP and AI techniques
Application goal AI technique Articles Count
[…]
Application goal AI technique Articles Count
[…]
Mult. AI techniques [105] 1
Premise selection DT [84, 91] 2
NN [96, 124, 139, 155, 174, 177, 178, 196, 209, 218, 237, 254] 12
RL [168, 241] 2
EA [188] 1
NLP [197, 217, 223] 3
Bayesian inference [221] 1
Mult. AI techniques [156, 173, 189, 111] 4
Custom [154, 214] 2
Proof search DT [144, 225] 2
KNN [122, 125] 2
NN [83, 92, 120, 123, 146, 165] 6
RL [94, 115, 130, 176, 191, 224, 250] 7
NLP [145] 1
Bayesian inference [136] 1
Mult. AI techniques [187, 261] 2
Proof mining NLP [183] 1
Data mining [143] 1
Proof synthesis NN [97, 99, 126, 175, 190, 228, 232, 257, 258] 9
RL [149, 207, 253] 3
NLP [131, 229, 236, 238] 4
Automaton learning [138] 1
Mult. AI techniques [208] 1
Symbolic RL [184] 1
Formula synthesis RL [148] 1
Formula classification NN [179] 1
Mult. AI techniques [180] 1
Tactic prediction NN [106] 1
RL [132] 1
Data mining [88] 1
Mult. AI techniques [172] 1
Lemma name generation NN [152] 1
Portfolio solving Mult. AI techniques [87] 1
Mult. Techniques NN [102, 129] 2
RL [167] 1
NLP [245] 1
Mult. AI techniques [266] 1
Entries grouped by SAT and AI techniques
Application goal AI technique Articles Count
[…]
Application goal AI technique Articles Count
[…]
NN [104, 203, 242, 159] 4
RL [101, 235] 2
Mult. AI techniques [112] 1
Custom [103] 1
Predict solvability DT [181] 1
NN [90, 121, 141, 219] 4
MaxSat NN [150, 170] 2
RL [98, 240] 2
EA [194, 200, 204, 213] 4
Custom [255] 1
SAT parameter selection RF [201] 1
Branching heuristics RL [127] 1
Mult. AI techniques [128] 1
Algorithm/solver selection RF [117] 1
SVM [133] 1
RL [142, 235] 2
Mult. AI techniques [85, 212] 2
Custom [60] 1
Instance selection NN [137, 158, 169, 262] 4
Clustering [86, 114] 2
Dependency analysis NN [259] 1
Meta analysis/performance prediction NN [151, 206, 234] 3
EA [147] 1
Generation NN [239, 110] 2
Mult. SAT tech NN [222] 1
EA [194] 1
Entries grouped by synthesis and AI techniques
Application goal AI technique Articles Count
[…]
Application goal AI technique Articles Count
[…]
RL [256, 260] 2
Loop invariant learning NN [192] 1
RL [267] 1
Clustering [244] 1
Custom [157] 1
Model repair Mult. AI techniques [80, 93, 215] 3
Models/Specification NN [210, 211] 2
RL [199] 1
NLP [116, 160, 202, 246] 4
Custom [247] 1
Annotations NLP [162] 1
Datasets
FM Articles Count
[…]
FM Articles Count
[…]
SMT [89] 1
TP [92, 97, 125, 126, 132, 138, 145, 146, 153, 180, 183, 209, 228, 241, 265] 15
MC [195] 1
Synthesis [80, 211, 246] 3